OpenStudy (anonymous):

Question 1.1. Which ordered pair makes the equation true? 2x – 5y = –10 (Points : 1) (5, 4) (–10, 4) (2, 5) (20, 4)

OpenStudy (anonymous):

plug them in and check

OpenStudy (anonymous):

\[2\times 5-5\times 4=10-20=-10\] got one!